# Sums of four fourth power of primes

### v2

@@ 3. Some Lemmas
$J(n)$ $\leq\frac{1+\varepsilon}{4\sqrt{2}}I^{*}U^{-2}L^{-1}\Bigg(\>\int_{U^{4}}^{16U^{4}}\>\frac{du}{u^{3/4}\log u}\Bigg)^{2}\Bigg(\>\int_{V^{4}}^{16V^{4}}\>\frac{dv}{v^{3/4}\log v}\Bigg)^{4}+O(V^{4}L^{-3B})$
$<\left(\frac{24}{5}\right)^{4}\frac{4}{\sqrt{2}}(1+\varepsilon)I^{*}V^{4}L^{-7}+O(V^{4}L^{-3B})$
$<11606V^{4}L^{-7},$
where we have used the estimate $I^{*}<7.73$. This proves (ii). For the proofs of (i) and (iii), we can refer to the methods of Brüdern [1] and Ren[2]. This proves Lemma 1.
@@
$y_{2}^{4}+\cdots+y_{4}^{4}-y_{5}^{4}-\cdots-y_{8}^{4}\equiv n(\text{mod}~p)$
where $1\leq y_{i}\leq p-1$ $(1\leq i\leq 8)$, then we get
$pK(n,p)=(p-1)^{7}+E,$
@@
$|E|\leqslant(3\sqrt{p}+1)^{5}(p-1)(3p+1),~~~\textit{if}~p\equiv 1({\mathrm{mod}}~4).$
@@
$pK(n,p)=\sum_{a=1}^{p}C^{3}(p,a)\overline{C^{4}(p,a)}e\left(-\frac{an}{p}\right)=(p-1)^{7}+E,$
@@
$E=\sum_{a=1}^{p-1}C^{3}(p,a)\overline{C^{4}(p,a)}e\biggl(-\frac{an}{p}\biggr).$
@@
Obiviously, $\sum_{a=1}^{p}|C(p,a)|^{2}$ can be expressed as $p$ times the number of solutions to equation $x^{4}\equiv y^{4}({\mathrm{mod~}}p)$, $1\leqslant x$, $y\leqslant p-1$. For $p\equiv 1({\mathrm{mod~}}4)$,
@@
$\sum_{a=1}^{p-1}|C^{3}(p,a)\overline{C^{4}(p,a)}|\leqslant(3\sqrt{p}+1)^{5}(p-1)(3p+1).$
@@
$x^{3}+y_{2}^{3}+\cdots+y_{4}^{3}-y_{5}^{3}-\cdots-y_{8}^{3}\equiv n\pmod{p^{i}}$
@@
$|E^{*}|\leq 3\sqrt{p}(3\sqrt{p}+1)^{5}(p-1)(3p+1),~~\textit{if}~p\equiv 1(\text{mod }4).$
It follows that $H(n,p)>0$ for all pp.
@@
$\prod_{p\leq x,p\equiv l({\mathrm{mod~}k})}\left(1-\frac{1}{p}\right)=\frac{e^{-\frac{1}{\varphi(k)}\left[\gamma+\ln\frac{\varphi(k)}{k}\right]}}{\ln^{\frac{1}{\varphi(k)}}x}\left\{1+O\left(e^{-c\ln^{\frac{3}{5}}x}\right)\right\}$
where $c$ is a positive absolute constant, $\gamma$ is the Euler’s constant, and the constant in $O$ is independent of $x$.
@@
$\sum_{N/9<\ell\leqslant N}r^{2}(\ell)<bV^{4}L^{-8},$
where $b=3702014554$.
Proof. Denote $\mathcal{B}$ as the set of all prime numbers greater than 13, and define
@@
$1+T_{p}(n,p)=\sum_{a=1}^{p}\frac{\overline{C(p,a)}|C(p,a)|^{6}}{(p-1)^{7}}e\Bigg(-\frac{an}{p}\Bigg)=p\frac{K(n,p)}{(p-1)^{7}}.$
@@
$\sum_{\begin{subarray}{c}m\sim U\\ m\equiv 0\mod d\end{subarray}}R(m)=\frac{\omega(d)}{d}X+E_{d}(n).$
Suppose $U_{0}\geqslant 1,\ V_{0}\geqslant 1,\ U_{0}V_{0}=D=N^{(1-\epsilon)/16}$ and $z=D^{1/2}$, then
@@
$\sum_{1\leqslant m\leqslant U}a_{m}\sum_{1\leqslant k\leqslant V}b_{k}E_{mk}\ll\sum_{1\leqslant d\leqslant D}\tau(d)E_{d}\ll V^{4}L^{-A}.$
By Lemma 5, we get
@@
(3.2) $\begin{split}&\mathfrak{S}_{1}(n)W(z)\\ =&\{1+T_{1}(n,2)+T_{1}(n,2^{2})+T_{1}(n,2^{3})+T_{1}(n,2^{4})\}\\ &\times(1+T_{1}(n,3))(1+T_{1}(n,5))(1+T_{1}(n,7))(1+T_{1}(n,11))\\ &\times\prod_{13\leqslant p<N^{(1-\varepsilon)/32}}(1+T_{1}(n,p))\biggl(1-\frac{K(n,p)}{H(n,p)}\biggr)\\ &\times\prod_{p\geqslant N^{(1-\varepsilon)/32}}(1+T_{1}(n,p))\\ =&\{1+T_{1}(n,2)+T_{1}(n,2^{2})+T_{1}(n,2^{3})+T_{1}(n,2^{4})\}\\ &\times(1+T_{1}(n,3))(1+T_{1}(n,5))(1+T_{1}(n,7))(1+T_{1}(n,11))\\ &\times\prod_{13\leqslant p<N^{(1-\varepsilon)/32}}\left(1-\frac{1}{p}\right)\left(1+\frac{E^{*}-E}{\left(p-1\right)^{8}}\right)\\ &\times\prod_{p\geqslant N^{(1-\varepsilon)/32}}\left(1+\frac{E^{*}}{p(p-1)^{7}}\right).\end{split}$
@@
$T_{1}(n,2)=\sum_{\overset{a=1}{(a,2)=1}}^{2}\frac{S(2,a)C^{3}(2,a)\overline{C^{4}(2,a)}}{2\varphi^{7}(2)}e\biggl(-\frac{an}{2}\biggr)=0.$
@@
$T_{1}(n,4)=\sum_{\overset{a=1}{(a,4)=1}}^{4}\frac{S(4,a)C^{3}(4,a)\overline{C^{4}(4,a)}}{4\varphi^{7}(4)}e\biggl(-\frac{an}{4}\biggr),$
@@
$\begin{split}T_{1}(n,4)&=\sum_{\overset{a=1}{(a,4)=1}}^{4}\frac{(2+2e(\frac{a}{4}))(2e(\frac{a}{4}))^{3}\overline{2e(\frac{a}{4})}^{4}}{4\varphi^{7}(4)}e\biggl(-\frac{a\times 0}{4}\biggr)\\ &=\frac{1}{2}\sum_{\overset{a=1}{(a,4)=1}}^{4}(1+e(-\frac{a}{4}))=\frac{1}{2}(2+2\cos{\frac{\pi}{2}})=1.\end{split}$
@@
$T_{1}(n,8)=\frac{1}{2}(4+2\cos{\frac{\pi}{4}}+2\cos{\frac{3\pi}{4}})=2$
@@
$T_{1}(n,16)=\frac{1}{2}(8+2\cos{\frac{\pi}{8}}+2\cos{\frac{3\pi}{8}}++2\cos{\frac{5\pi}{8}}+2\cos{\frac{7\pi}{8}})=4.$
@@
$T_{1}(n,3)=\sum_{\overset{a=1}{(a,3)=1}}^{3}\frac{S(3,a)C^{3}(3,a)\overline{C^{4}(3,a)}}{3\varphi^{7}(3)}e\biggl(-\frac{an}{3}\biggr),$
@@
$S(3,a)=\sum_{m=1}^{3}e\biggl(\frac{am^{4}}{3}\biggr)=1+2e(\frac{a}{3}),C(4,a)=\sum_{(m,4)=1}^{4}e\biggl(\frac{am^{4}}{4}\biggr)=2e(\frac{a}{3}).$
@@
$\begin{split}T_{1}(n,3)&=\sum_{\overset{a=1}{(a,3)=1}}^{3}\frac{(1+2e(\frac{a}{3}))(2e(\frac{a}{3}))^{3}\overline{(2e(\frac{a}{3}))^{4}}}{3\varphi^{7}(3)}e\biggl(-\frac{a\times 0}{3}\biggr)\\ &=\frac{1}{3}(4+e(-\frac{1}{3})+e(-\frac{2}{3}))=\frac{1}{3}(4+2\cos{\frac{2\pi}{3}})=1.\end{split}$
Likewise, by a series of calculations we have $T_{1}(n,5)=\frac{1}{5}(16+2\cos{\frac{2\pi}{5}}+2\cos{\frac{4\pi}{5}})=\frac{1}{5}(16+\frac{\sqrt{5}-1}{2}-\frac{\sqrt{5}+1}{2})=3$, $T_{1}(n,7)=\frac{8}{729}$ and $T_{1}(n,11)=\frac{324}{171875}.$
@@
$\prod_{13\leqslant p<N^{(1-\varepsilon)/32}}\left(1+\frac{E^{*}-E}{\left(p-1\right)^{8}}\right),$

### v5

@@ 3. Some Lemmas
$J(n)$ $\leq\frac{1+\varepsilon}{4\sqrt{2}}I^{*}U^{-2}L^{-1}\Bigg(\>\int_{U^{4}}^{16U^{4}}\>\frac{du}{u^{3/4}\log u}\Bigg)^{6}\Bigg(\>\int_{V^{4}}^{16V^{4}}\>\frac{dv}{v^{3/4}\log v}\Bigg)^{8}+O(U^{4}V^{8}L^{-5B})$
$<\left(\frac{32}{7}\right)^{8}\frac{4^{5}}{\sqrt{2}}(1+\varepsilon)I^{*}U^{4}V^{8}L^{-15}+O(U^{4}V^{8}L^{-5B})$
$<4888799.222U^{4}V^{8}L^{-15},$
where we have used the estimate $I^{*}<7.73$. This proves (ii). Now we prove (iii), let
$v_{d}(n)=\sum_{\begin{subarray}{c}m\sim U\\ m\equiv 0({\mathrm{mod}}~d)\end{subarray}}R(m);f_{d}(\alpha)=\sum_{U<x\leq 2U\atop x\equiv 0\mod d}e(\alpha x^{4});$
$g(\alpha)=\sum_{U<p\leq 2U}e(\alpha p^{4});h(\alpha)=\sum_{V<p\leq 2V}e(\alpha p^{4})$
and then define
$F(\alpha)=\sum_{d<D}\eta_{d}f_{d}(\alpha),$
$v_{d}(n,\mathfrak{B})=\int_{\mathfrak{B}}f_{d}(\alpha)\overline{g(\alpha)}|g(\alpha)|^{4}|h(\alpha)|^{8}e(-\alpha n)d\alpha.$
Naturally, $v_{d}(n,[0,1])=v_{d}(n)$, therefore
(3.4) $\begin{split}|\sum_{d\leqslant D}\eta_{d}E_{d}(n)|&\leq\sum_{d\leqslant D}|\eta_{d}||v_{d}(n)-\frac{\mathfrak{S}_{d}(n)}{d}J(n)|\\ &\leq\sum_{d\leqslant D}|\eta_{d}||v_{d}(n,\mathfrak{M})-\frac{\mathfrak{S}_{d}(n)}{d}J(n)|+\sum_{d\leqslant D}|\eta_{d}||v_{d}(n,\mathfrak{m})|.\end{split}$
Next, we calculate the second part on the right side of the inequality
(3.5) $\begin{split}\sum_{d\leqslant D}|\eta_{d}||v_{d}(n,\mathfrak{m})|&=\int_{\mathfrak{m}}|F(\alpha)||g(\alpha)|^{7}|h(\alpha)|^{8}e(-\alpha n)d\alpha\\ &\leq\bigg(\int_{\mathfrak{m}}|F(\alpha)|^{2}|g(\alpha)|^{2}|h(\alpha)|^{4}d\alpha\bigg)^{1/2}\bigg(\int_{\mathfrak{m}}|g(\alpha)|^{12}|h(\alpha)|^{12}d\alpha\bigg)^{1/2}.\end{split}$
Now we calculate the upper bound of $F(\alpha)$. Regarding the above $\mathfrak{M}^{\prime}(q,a)$, $1\leq a\leq q\leq U^{2}$, we write $F(\alpha)$ as
$F(\alpha)=\sum_{d\leq D}\eta_{d}\sum_{U/d<y\leq 2U/d}e(\alpha d^{4}y^{4}).$
By Dirichlet’ s theorem on diophantine approximation, there are coprime integers $b$, $r$ with $r\leq 16P^{3}d^{-3}$, $\left|d^{4}\alpha-\frac{b}{r}\right|\leq\frac{1}{16}r^{-1}d^{3}P^{-3}$. By Weyl’ s inequality, when $r>U/d$
$\sum_{U/d<y\leq 2U/d}e(\alpha d^{4}y^{4})\ll\bigg(\frac{U}{d}\bigg)^{7/8},$
and when $r\leq U/d$
(3.6) $\sum_{U/d<y\leq 2U/d}e(\alpha d^{4}y^{4})\ll r^{-\frac{1}{4}}\frac{U}{d}\bigg(1+\bigg(\frac{U}{d}\bigg)^{4}\bigg|\alpha d^{4}-\frac{b}{r}\bigg|\bigg)^{-\frac{1}{4}}+\bigg(\frac{P}{d}\bigg)^{\frac{1}{2}+\epsilon}.$
Furthermore, when
$r\leq(U/d)^{7/8};\quad\left|\alpha d^{4}-\frac{b}{r}\right|\leq\frac{1}{r}\left(\frac{d}{U}\right)^{\frac{7}{2}},$
it can also achieve a result of (3.6). Hence
$F(\alpha)\ll U^{7/8+\epsilon}D^{1/8}+U\sum_{d\in\mathcal{D}}d^{-1}r^{-\frac{1}{4}}\biggl(1+\biggl(\frac{U}{d}\biggr)^{4}\biggl|\alpha d^{4}-\frac{b}{r}\biggr|\biggr)^{-\frac{1}{4}},$
where $\mathcal{D}$ represents the set of d that satisfies the condition. Compare the conditions of $q,a$ and $b,r$, we have
$\bigg|\frac{b}{r}-\frac{ad^{4}}{q}\bigg|\leq\frac{1}{r}\bigg(\frac{d}{U}\bigg)^{\frac{7}{2}}+\frac{d^{4}}{qU^{2}}$
i.e.
$\bigg|bq-ad^{4}r\bigg|\leq qd^{\frac{7}{2}}U^{-\frac{7}{2}}+rU^{-2}D^{4}\ll 1$
where $D=N^{\xi}$, $0<\xi<9/25$, $U$ is large enough. Therefore $bq=ad^{4}r$, then $r=q/(q,d^{4})$, by the trivial bound $(q,d^{4})\leq(q,d)^{4}$
$\sum_{d\in\mathcal{D}}d^{-1}r^{-\frac{1}{4}}\left(1+\left(\frac{U}{a}\right)^{4}\left|\alpha d^{4}-\frac{b}{r}\right|\right)^{-\frac{1}{4}}\leq q^{-\frac{1}{4}}\left(1+P^{4}\left|\alpha-\frac{a}{q}\right|\right)^{-\frac{1}{4}}\sum_{d\leq D}\frac{(q,d)}{d}.$
Thus
$F(\alpha)\ll U^{\frac{7}{8}+\epsilon}D^{\frac{1}{8}}+q^{\epsilon-\frac{1}{4}}P(\log P)\left(1+P^{4}\bigg|\alpha-\frac{a}{q}\bigg|\right)^{-\frac{1}{4}}.$
Let $\delta>0$ be so small that $D^{\frac{1}{8}}U^{\frac{7}{8}}\leq U^{\frac{23}{25}-2\delta}$, this is always possible because $D<U^{\frac{9}{25}}$. Let $\mathfrak{N}(q,a)$ denote the interval $|q\alpha-a|\leqslant U^{5\delta-\frac{92}{25}}$, and $\mathfrak{N}$ be the union of all $\mathfrak{N}(q,a)$ with $1\leqslant a\leqslant q\leqslant P^{\frac{8}{25}+5\delta}$, $(a,q)=1$. By the upper bound of $F(\alpha)$ obtained from the above, for $|F(\alpha)>U^{\frac{23}{25}}|$, then $\alpha\in\mathfrak{N}$. Defining $\Phi$ on $\mathfrak{N}$ as
$\Phi(\alpha)=q^{\epsilon-\frac{1}{4}}\Bigg(1+U^{4}\Bigg|\alpha-\frac{a}{q}\Bigg|\Bigg)^{-1/4}$
Then we have
(3.7) $\int_{\mathfrak{m}}|Fgh^{2}|^{2}d\alpha\ll U^{\frac{46}{25}-2\delta}\int_{0}^{1}|gh^{2}|^{2}d\alpha+U^{2}(\log U)^{2}\int_{\mathfrak{N}\cap\mathfrak{m}}|\Phi gh^{2}|^{2}d\alpha.$
By Hölder’ s inequality
(3.8) $\int_{\mathfrak{N}\cap\mathfrak{m}}|\Phi gh^{2}|^{2}d\alpha\leq\left(\int_{\mathfrak{N}\cap\mathfrak{m}}|\Phi|^{12}d\alpha\right)^{\frac{1}{6}}U_{1}^{\frac{1}{2}}U_{2}^{\frac{1}{3}}$
where
$U_{1}=\int_{0}^{1}|g(\alpha)|^{2}|h(\alpha)|^{4}d\alpha;\quad U_{2}=\int_{0}^{1}|g(\alpha)|^{3}|h(\alpha)|^{6}d\alpha.$
By simple calculations we obtain that
$\int_{\mathfrak{N}\cap\mathfrak{m}}|\Phi(\alpha)|^{12}d\alpha\ll U^{-4}L^{-B}.$
Then let $\mathfrak{m}_{1}(q,a)$ denote the interval $|q\alpha-a|\leqslant U^{\frac{12}{5}}$, $\mathfrak{m}_{1}$ be the union of all $\mathfrak{m}_{1}(q,a)$ with $1\leqslant a\leqslant q\leqslant U^{\frac{4}{5}}$, $(a,q)=1$, and $\mathfrak{m}_{2}(q,a)$ denote the interval $|q\alpha-a|\leqslant V^{\frac{12}{5}}$, $\mathfrak{m}_{2}$ be the union of all $\mathfrak{m}_{2}(q,a)$ with $1\leqslant a\leqslant q\leqslant V^{\frac{4}{5}}$, $(a,q)=1$. By Theorem 4.1, Lemma 6.3 of Vaughan[7], and Theorem 2 of Vaughan[11], we have
$g(\alpha)\ll q^{-\frac{1}{4}}U\bigg(1+U^{4}\bigg|\alpha-\frac{a}{q}\bigg|\bigg)^{-1}+U^{4/5+\epsilon}$
where $\alpha\in\mathfrak{m}_{1}(q,a)$;
$h(\alpha)\ll q^{-\frac{1}{4}}V\bigg(1+V^{4}\bigg|\alpha-\frac{a}{q}\bigg|\bigg)^{-1}+V^{4/5+\epsilon}$
where $\alpha\in\mathfrak{m}_{2}(q,a).$ Moreover, $|g(\alpha)|>U^{4/5+2\epsilon}$ implies $\alpha\in{\mathfrak{m}_{1}}({\mathrm{mod}}~1)$, $|h(\alpha)|>V^{4/5+2\epsilon}$ implies $\alpha\in{\mathfrak{m}_{2}}({\mathrm{mod}}~1)$.
(3.9) $g(\alpha)\ll q^{-1/4}U+U^{1/2}\ll q^{-1/4}U$
where $\alpha\in{\mathfrak{m}_{1}}({\mathrm{mod}}~1)$;
(3.10) $h(\alpha)\ll q^{-1/4}V+V^{1/2}\ll q^{-1/4}V$
where $\alpha\in{\mathfrak{m}_{2}}({\mathrm{mod}}~1).$ Then define $\Psi(\alpha)$ on $\mathfrak{m}_{1}$, $\Psi^{*}(\alpha)$ on $\mathfrak{m}_{2}$ as
(3.11) $\Psi(\alpha)=q^{-\frac{1}{4}}\left(1+U^{4}\bigg|\alpha-\frac{a}{q}\bigg|\right)^{-1};\Psi^{*}(\alpha)=q^{-\frac{1}{4}}\left(1+V^{4}\bigg|\alpha-\frac{a}{q}\bigg|\right)^{-1},$
Subsequently
$\int_{0}^{1}|g(\alpha)|^{3}|h(\alpha)|^{6}d\alpha\ll$ $\>U^{\frac{4}{5}}\int_{0}^{1}|g(\alpha)|^{2}|h(\alpha)|^{6}d\alpha+U\int_{\mathfrak{m}_{1}}|\Psi(\alpha)|^{2}|g(\alpha)|^{2}|h(\alpha)|^{6}d\alpha$
$\ll$ $\>U^{\frac{4}{5}}V^{\frac{8}{5}}\int_{0}^{1}|g(\alpha)|^{2}|h(\alpha)|^{4}d\alpha+U^{\frac{4}{5}}V^{2}\int_{\mathfrak{m}_{2}}|\Psi^{*}(\alpha)|^{2}|g(\alpha)|^{2}|h(\alpha)|^{4}d\alpha$
$+UV^{\frac{8}{5}}\int_{\mathfrak{m}_{1}}|\Psi(\alpha)|^{2}|g(\alpha)|^{2}|h(\alpha)|^{4}d\alpha$
$+UV^{2}\int_{\mathfrak{m}_{1}\cap\mathfrak{m}_{2}}|\Psi(\alpha)|^{2}|\Psi^{*}(\alpha)|^{2}|g(\alpha)|^{2}|h(\alpha)|^{4}d\alpha.$
By (3.9), (3.10), (3.11) we can obtain that
(3.12) $U_{2}=\int_{0}^{1}|g(\alpha)|^{3}|h(\alpha)|^{6}d\alpha\ll U^{5+13/40}.$
Since the upper bound of $U_{1}$ is known, we can combine(3.7), (3.8), (3.12) to obtain that
$\int_{\mathfrak{m}}|Fgh^{2}|^{2}d\alpha\ll U^{4+193/200}L^{-B/6}.$
Similar to the method mentioned above, we can obtain
$\int_{\mathfrak{m}}|g(\alpha)|^{12}|h(\alpha)|^{12}d\alpha\ll U^{16+29/40}$
Naturally, by (3.5)
$\sum_{d\leqslant D}|\eta_{d}||v_{d}(n,\mathfrak{m})|\ll U^{11}L^{-B/12}$
Now we calculate the first term on the right side of the inequality (3.4). For $\alpha=a/q+\beta\in\mathfrak{M}$, we define
$f_{d}^{*}(\alpha)=\frac{S(q,ad^{4})}{qd}\Phi(\lambda,U),\quad g^{*}(\alpha)=\frac{C(q,a)}{\varphi(q)}\Psi(\lambda,U),\quad h^{*}(\alpha)=\frac{C(q,a)}{\varphi(q)}\Psi(\lambda,V).$
By Lemma 7.15 of Hua[12]
$g(\alpha)-g^{*}(\alpha)\ll U\exp(-c_{1}\sqrt{L}),\quad h(\alpha)-h^{*}(\alpha)\ll V\exp(-c_{1}\sqrt{L})$
where $c_{1}>0$ is a absolute constant.By Theorem 4.1 of Vaughan[11]
$f_{d}(\alpha)-f_{d}^{*}(\alpha)\ll q^{1/2+\epsilon}(1+U^{4}|\lambda|)\ll L^{2B}.$
Then for $\alpha\in\mathfrak{M}$, by trivial estimate $\left|S(q,ad^{3})\right|\leq q$, we have
$f_{d}\overline{g}|g|^{6}|h|^{8}-f_{d}^{*}\overline{g^{*}}|g^{*}|^{6}|h^{*}|^{8}\ll d^{-1}U^{8}V^{8}L^{-13B},$
and
(3.13) $\sum_{d\leq D}|\eta_{d}||\nu_{d}(n,\mathfrak{M})-\nu_{d}^{*}(n)|\ll U^{4}V^{8}L^{-A},$
where
$\nu_{d}^{*}(n)=\int_{\mathfrak{M}}f_{d}^{*}\overline{{g^{*}}}|g^{*}|^{6}|h^{*}|^{8}e(-n\alpha)d\alpha,$
and we can find that
(3.14) $\nu_{d}^{*}(n)=\frac{1}{d}J(n)\sum_{q\leq L^{B}}T_{d}(n,q).$
Then by $|C(q,a)|\ll q^{3/4+\epsilon}$
$|T_{d}(n,q)|\leq\sum_{a=1}^{q}\frac{|S(q,ad^{4})||C(q,a)|^{15}}{q\varphi^{15}(q)}\ll\sum_{a=1}^{q}\frac{q^{49/4+\epsilon}}{q\varphi^{15}(q)}\ll q^{-2}.$
Hence the series
$\mathfrak{S}_{d}(n)=\sum_{q=1}^{\infty}T_{d}(n,q)$
converges absolutely, this proves (i), and
$\sum_{q\leq L^{B}}T_{d}(n,q)=\mathfrak{S}_{d}(n)+O(L^{-B}).$
By (3.14)
$\nu_{d}^{*}(n)=\frac{\mathfrak{S}_{d}(n)}{d}J(n)+O\left(\frac{U^{4}V^{8}}{dL^{B}}\right),$
combining with (3.13), we can obtain
$\sum_{d\leqslant D}|\eta_{d}||v_{d}(n,\mathfrak{M})-\frac{\mathfrak{S}_{d}(n)}{d}J(n)|\ll U^{4}V^{8}L^{-A}.$
This proves (iii), which also proves Lemma 2.
@@
$y_{2}^{4}+\cdots+y_{8}^{4}-y_{9}^{4}-\cdots-y_{16}^{4}\equiv n(\text{mod}~p)$
where $1\leq y_{i}\leq p-1$ $(2\leq i\leq 16)$, then we get
$pK(n,p)=(p-1)^{15}+E,$
@@
$|E|\leqslant(3\sqrt{p}+1)^{{13}}(p-1)(3p+1),~~~\textit{if}~p\equiv 1({\mathrm{mod}}~4).$
@@
$pK(n,p)=\sum_{a=1}^{p}C^{7}(p,a)\overline{C^{8}(p,a)}e\left(-\frac{an}{p}\right)=(p-1)^{15}+E,$
@@
$E=\sum_{a=1}^{p-1}C^{7}(p,a)\overline{C^{8}(p,a)}e\biggl(-\frac{an}{p}\biggr).$
@@
Obviously, $\sum_{a=1}^{p}|C(p,a)|^{2}$ can be expressed as $p$ times the number of solutions to equation $x^{4}\equiv y^{4}({\mathrm{mod~}}p)$, $1\leqslant x$, $y\leqslant p-1$. For $p\equiv 1({\mathrm{mod~}}4)$,
@@
$\sum_{a=1}^{p-1}|C^{7}(p,a)\overline{C^{8}(p,a)}|\leqslant(3\sqrt{p}+1)^{13}(p-1)(3p+1).$
@@
$x^{4}+y_{2}^{4}+\cdots+y_{8}^{4}-y_{9}^{4}-\cdots-y_{16}^{4}\equiv n\pmod{p^{i}}$
@@
$|E^{*}|\leq 3\sqrt{p}(3\sqrt{p}+1)^{13}(p-1)(3p+1),~~\textit{if}~p\equiv 1(\text{mod }4).$
It also follows that $H(n,p)>0$ for $p>17$.
@@
$\prod_{\begin{subarray}{c}p\leq x\\ p\equiv l({\mathrm{mod~}k})\end{subarray}}\left(1-\frac{1}{p}\right)=\frac{e^{-\frac{1}{\varphi(k)}\left[\gamma+\ln\frac{\varphi(k)}{k}\right]}}{\ln^{\frac{1}{\varphi(k)}}x}\left\{1+O\left(e^{-c\ln^{\frac{3}{5}}x}\right)\right\}$
where $c$ is a positive absolute constant, $\gamma$ is the Euler’s constant, and the constant in $O$ is independent of $x$. Proof. See [15], corollary of Theorem 429.
@@
$\sum_{N/9<\ell\leqslant N}r^{2}(\ell)<bU^{2}V^{8}L^{-14},$
where $b=80947432211.141$.
Proof. Denote $\mathcal{B}$ as the set of all prime numbers greater than 17, and define
@@
$1+T_{p}(n,p)=\sum_{a=1}^{p}\frac{\overline{C(p,a)}|C(p,a)|^{14}}{(p-1)^{15}}e\Bigg(-\frac{an}{p}\Bigg)=p\frac{K(n,p)}{(p-1)^{15}}.$
@@
$\sum_{\begin{subarray}{c}m\sim U\\ m\equiv 0(\mod d)\end{subarray}}R(m)=\frac{\omega(d)}{d}X+E_{d}(n).$
Suppose $U_{0}\geqslant 1,\ V_{0}\geqslant 1,\ U_{0}V_{0}=D=N^{9(1-\epsilon)/100}$ and $z=D^{1/2}$, then
@@
$\sum_{1\leqslant m\leqslant U}a_{m}\sum_{1\leqslant k\leqslant V}b_{k}E_{mk}\ll\sum_{1\leqslant d\leqslant D}\tau(d)E_{d}\ll U^{4}V^{8}L^{-A}.$
By the upper bound in Lemma 6, we get
@@
(3.15) $\begin{split}&\mathfrak{S}_{1}(n)W(z)\\ =&\{1+T_{1}(n,2)+T_{1}(n,2^{2})+T_{1}(n,2^{3})+T_{1}(n,2^{4})\}\\ &\times(1+T_{1}(n,3))(1+T_{1}(n,5))(1+T_{1}(n,7))(1+T_{1}(n,11))(1+T_{1}(n,13))\\ &\times\prod_{17\leqslant p<N^{9(1-\varepsilon)/200}}(1+T_{1}(n,p))\biggl(1-\frac{K(n,p)}{H(n,p)}\biggr)\\ &\times\prod_{p\geqslant N^{9(1-\varepsilon)/200}}(1+T_{1}(n,p))\\ =&\{1+T_{1}(n,2)+T_{1}(n,2^{2})+T_{1}(n,2^{3})+T_{1}(n,2^{4})\}\\ &\times(1+T_{1}(n,3))(1+T_{1}(n,5))(1+T_{1}(n,7))(1+T_{1}(n,11))(1+T_{1}(n,13))\\ &\times\prod_{17\leqslant p<N^{(1-\varepsilon)/200}}\left(1-\frac{1}{p}\right)\left(1+\frac{E^{*}-E}{\left(p-1\right)^{16}}\right)\\ &\times\prod_{p\geqslant N^{(1-\varepsilon)/200}}\left(1+\frac{E^{*}}{p(p-1)^{15}}\right).\end{split}$
@@
$T_{1}(n,2)=\sum_{\overset{a=1}{(a,2)=1}}^{2}\frac{S(2,a)C^{7}(2,a)\overline{C^{8}(2,a)}}{2\varphi^{16}(2)}e\biggl(-\frac{an}{2}\biggr)=0.$
@@
$T_{1}(n,4)=\sum_{\overset{a=1}{(a,4)=1}}^{4}\frac{S(4,a)C^{7}(4,a)\overline{C^{8}(4,a)}}{4\varphi^{16}(4)}e\biggl(-\frac{an}{4}\biggr),$
@@
$\begin{split}T_{1}(n,4)&=\sum_{\overset{a=1}{(a,4)=1}}^{4}\frac{(2+2e(\frac{a}{4}))(2e(\frac{a}{4}))^{7}\overline{2e(\frac{a}{4})}^{8}}{4\varphi^{13}(4)}e\biggl(-\frac{a\times 0}{4}\biggr)\\ &=\frac{S(4,1)C^{7}(4,1)+S(4,3)C^{7}(4,3)}{2^{15}}=1.\end{split}$
@@
$T_{1}(n,8)=\sum_{\overset{a=1}{(a,8)=1}}^{8}\frac{1}{2}\bigg(1+e\bigg(\frac{7a}{8}\bigg)\bigg)=2$
@@
$T_{1}(n,16)=\sum_{\overset{a=1}{(a,16)=1}}^{16}\frac{1}{2}\bigg(1+e\bigg(\frac{15a}{16}\bigg)\bigg)=4.$
@@
$T_{1}(n,3)=\sum_{\overset{a=1}{(a,3)=1}}^{3}\frac{S(3,a)C^{7}(3,a)\overline{C^{8}(3,a)}}{3\varphi^{15}(3)}e\biggl(-\frac{an}{3}\biggr),$
@@
$S(3,a)=\sum_{m=1}^{3}e\biggl(\frac{am^{4}}{3}\biggr)=1+2e\bigg(\frac{a}{3}\bigg),C(4,a)=\sum_{(m,4)=1}^{4}e\biggl(\frac{am^{4}}{4}\biggr)=2e\bigg(\frac{a}{3}\bigg).$
@@
$\begin{split}T_{1}(n,3)&=\sum_{\overset{a=1}{(a,3)=1}}^{3}\frac{(1+2e(\frac{a}{3}))(2e(\frac{a}{3}))^{7}\overline{(2e(\frac{a}{3}))^{8}}}{3\varphi^{15}(3)}e\biggl(-\frac{a\times 0}{3}\biggr)=1.\end{split}$
Likewise, by a series of calculations we have $T_{1}(n,5)=3$, $T_{1}(n,7)=\frac{128}{3^{14}}$, $T_{1}(n,11)=\frac{2187}{6103515625}$ and $T_{1}(n,13)\leq 0.015$ .
@@
$\prod_{17\leqslant p<N^{9(1-\varepsilon)/200}}\left(1+\frac{E^{*}-E}{\left(p-1\right)^{16}}\right),$